OpenStudy (anonymous):

A rectangle has one corner in Quadrant 1 on the graph of an ellipse whose equation is 9x^2+16y^2=144, another at the origin, a third on the positive y-axis, and the fourth on the positive x-axis. Express the area of the rectangle as a function of x.

OpenStudy (tkhunny):

What's your plan? It's a rectangle with opposite corners at the Origin and on the ellipse.

OpenStudy (anonymous):

The answers are, A. A(x)=Square root of .5625-9x^2 B. A(x)=x square root of 9-16x^2 C. A(x)=x square root of 16-9x^2 D. A(x)= x square root of 9-.5625x^2

OpenStudy (tkhunny):

Super. What's your plan? It's a rectangle with opposite corners at the Origin and on the ellipse.

OpenStudy (anonymous):

What does that mean?